b=24-2s. Solve for s
show steps plz

Answer by tommyt3rd(5050)

b=24-2s
b-24=24-24-2s
b-24=-2s
(b-24)/(-2)=-2s/(-2)
(b-24)/(-2)=s
s=-(b-24)/2